OpenStudy (anonymous):

12<2^(m/n)<13 ..... what are the smallest possible integers for m and n?

OpenStudy (anonymous):

its a challenging quesion...

OpenStudy (anonymous):

positive integers?

OpenStudy (anonymous):

n=3 and m=11 ?

OpenStudy (anonymous):

@aceace

OpenStudy (shubhamsrg):

let m/n =q => 12 < 2^q < 13 clearly,q will be between 3 and 4 2^q = 12 shall give our ans taking log to base 10, q(0.3) = 2(0.3) + (0.5) q = 11/3 so i go with @mukushla

OpenStudy (shubhamsrg):

wolfram tells me 3.485 approx => 697/200

OpenStudy (anonymous):

can you please explain the log part...

OpenStudy (shubhamsrg):

log(2) and log(3) to base 10 are basic ones.. log2 = 0.3010 and log 3 = 0.5 approx..

OpenStudy (anonymous):

\[12<2^{m/n}<13 \\ take \ \ \log \\ \log 12<\frac{m}{n} \log 2<\log 13 \\ \frac{\log 13}{\log 2}<\frac{m}{n} <\frac{\log 13}{\log 2}\\ 3.5<\frac{m}{n}<3.7\]

OpenStudy (anonymous):

simply u can see m=11 and n=3 is your answer

OpenStudy (anonymous):

how do you see that those integers are the answer?

OpenStudy (anonymous):

how do you see that those integers are the answer?

OpenStudy (anonymous):

how do you see that those integers are the answer?

OpenStudy (anonymous):

how do you see that those integers are the answer?

OpenStudy (anonymous):

how do you see that those integers are the answer?

OpenStudy (anonymous):

how do you see that those integers are the answer?

OpenStudy (anonymous):

how do you see that those integers are the answer?

OpenStudy (shubhamsrg):

note that inequality sign didnt reverse since we were taking base >1

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

i see.. thanks